Nicole:

Find the value of y so that the line passing through (6, y) and (7, 5) has a slope of -3.

Nicole:

@dude

dude:

So use the slope from 2 points formula \(\dfrac{y_2-y_1}{x_2-x_1}=m\) Substitute what you have (You try)

Nicole:

5-y / 7-6 5-y / 1

Nicole:

\[\frac{ 5-y }{ 7-6 }=-3\]

dude:

Good You had already started to solve it but yeah you would get to \({5-y}=-3\)

Nicole:

y=8

Nicole:

thats it right?

dude:

Ye